OpenStudy (anonymous):

passing through (3,-1) and parallel to y=2x-3

OpenStudy (anonymous):

Nevermind that, sorry it's y = 2x - 7

OpenStudy (anonymous):

Look go here: https://www.desmos.com/calculator And put y=2x-3 and y=2x-7 in. You will see that y=2x-7 lies on 3,-1

OpenStudy (anonymous):

slope is the coeeffficient of x i.e.2 and given point s(x1,y1)=(3,-1) the requird eqn, (y-y1)=m(x-x1) (y-(-1))=2(x-3) (y+1)=2x-6 y-2x=-6-1 y-2x=-7